# CustomCrafter: Customized Video Generation with Preserving Motion and Concept Composition Abilities

### v1

@@ Introduction
In this section, we construct ablation studies to validate the specific roles and effectiveness of each component that our proposed. As shown in  Table 2, we choose Custom Diffusion as the baseline to present the quantitative results of our designed Spatial Subject Learning Module and Dynamic Weighted Video Sampling Strategy, respectively. It can be observed that using only our Spatial Subject Learning Module achieves better results on the CLIP-I and DINO-I metrics, which measure the similarity of the subject’s appearance. This suggests that compared to previous work, our method is more capable of capturing the details of the subject. The Dynamic Weighted Video Sampling Strategy, due to modifications in the process, may result in a slight impairment of the subject’s appearance similarity. However, the motions can be significantly improved, substantially enhancing the congruence between the generated video and the text, as measured by the CLIP-T metrics. In addition, we use visualization results to demonstrate the effectiveness of our method. As shown in  Figure 7, when we only update the parameters of cross-attention, we find that the model’s ability to combine concepts is poor and it cannot generate a simple concept combination of pandas and bamboo that matches the prompt. However, when we use our Spatial Subject Learning Module module to update both cross-attention and self-attention, and do not use the Dynamic Weighted Video Sampling Strategy, the generated video’s ability to combine concepts is improved, but the VDM cannot generate fluent motions that match the text prompt. After adopting our sampling strategy, the generated video has almost no significant loss in subject appearance similarity, but the naturalness of the motion has greatly improved.
@@ Conclusion
In this paper, we introduce our CustomCrafter, a novel framework for customized video generation. This approach does not require additional video to repair motion generation ability. We first designed a Spatial Subject Learning Module, which updates the Spatial Attention to complete the learning of the subject’s appearance features. Simultaneously, we proposed a Dynamic Weighted Video Generation, which improves the model’s inference process to restore the motion generation capability of VDMs. Through qualitative and quantitative experiments, we have demonstrated that our method is better than existing methods, preserving the abilities of VDMs to combine concepts and generate motion.

### v2

@@ Introduction
In this section, we construct ablation studies to validate the specific roles and effectiveness of each component that our proposed. As shown in  Table 2, we choose Custom Diffusion as the baseline to present the quantitative results of our designed Spatial Subject Learning Module and Dynamic Weighted Video Sampling Strategy, respectively. It can be observed that using our Spatial Subject Learning Module achieves better results on the CLIP-I and DINO-I metrics, which measure the subject fidelity. This suggests that compared to previous work, our method is more capable of capturing the details of the subject. The Dynamic Weighted Video Sampling Strategy, due to modifications in the process, may result in a slight impairment of the subject’s appearance similarity. However, the motions can be significantly improved, substantially enhancing the text alignment. In addition, we use visualization results to demonstrate the effectiveness of our method. As shown in  Figure 7, when we only update the parameters of cross-attention, we find that the model’s ability to combine concepts is poor and it cannot generate a simple concept combination of pandas and bamboo that matches the prompt. However, when we use our Spatial Subject Learning Module module to update both cross-attention and self-attention, and do not use the Dynamic Weighted Video Sampling Strategy, the generated video’s ability to combine concepts is improved, but the VDM cannot generate fluent motions that match the text prompt. After adopting our sampling strategy, the generated video has almost no significant loss in subject fidelity, but the naturalness of the motion has greatly improved.
@@ Conclusion
In this paper, we introduce our CustomCrafter, a novel framework for customized video generation. This approach does not require additional video to repair motion generation ability. We first designed a Spatial Subject Learning Module, which updates the Spatial Attention for subject learning. Simultaneously, we proposed a Dynamic Weighted Video Sampling Strategy, which improves the model’s inference process to restore the motion generation capability of VDMs. Through experiments, we have demonstrated that our method is better than existing methods.
Acknowledgements
This work is supported in part by National Science Foundation for Distinguished Young Scholars under Grant 62225605, Project 12326608 supported by NSFC, Natural Science Foundation of Shanghai under Grant 24ZR1425600, Zhejiang Provincial Natural Science Foundation of China under Grant LD24F020016, “Pioneer” and “Leading Goose” R&D Program of Zhejiang (No. 2024C01020), as well as the Ningbo Science and Technology Innovation Project (No.2024Z294), and sponsored by CCF-Tencent Rhino-Bird Open Research Fund and Research Fund of ARC Lab, Tencent PCG.